OpenStudy (anonymous):

how do u solve for r? 2112=42(r)^8

OpenStudy (mathmale):

\[2112=42(r)^8\]

OpenStudy (mathmale):

... is easier to read. What would be your first step in solving this for r?

OpenStudy (anonymous):

divide both sides by 42

OpenStudy (anonymous):

@mathmale

OpenStudy (mathmale):

Good thinking. Yes, that's what I'd do. Start with \[2112=42(r)^8\]

OpenStudy (mathmale):

and divide both sides by 42. Your result?

OpenStudy (anonymous):

50.28=r^8

OpenStudy (mathmale):

Great. Let me ask you this: Have you already studied and used logarithms?

OpenStudy (anonymous):

use but i dont know how to apply that here

OpenStudy (anonymous):

yes*

OpenStudy (mathmale):

Have you learned how to raise bases such as 10 to fractional powers, such as \[10^{1/2}\]?

OpenStudy (anonymous):

i think but i dont quite remember

OpenStudy (mathmale):

Good. Now, back to your first result: 50.28=r^8. We want to solve that for r. One way to do that would be to take the 8th root of both sides of the equation:\[(50.28)^{1/8}=(r^8)^{1/8}\]

OpenStudy (mathmale):

Are you able to do these operations on your calculator? Actually, you don't need a calculator for the right side, since \[(r ^{8})^{1/8}=r ^{8/8}=r^1=r\]

OpenStudy (mathmale):

What kind of calculator have you?

OpenStudy (anonymous):

casio

OpenStudy (anonymous):

scientific

OpenStudy (mathmale):

Know how to evaluate 3^7 on your calculator?

OpenStudy (anonymous):

yes

OpenStudy (mathmale):

Know how to evaluate 3^(1/2) on your calculator? If so, find the approx value of 3^(1/2).

OpenStudy (anonymous):

1.73

OpenStudy (mathmale):

Hint: Be sure to enclose the exponent 1/2 in parentheses: 3^(1/2). Yes, that's correct. Using this approach, can you evaluate \[50.28^{1/2}?\]

OpenStudy (anonymous):

7.09

OpenStudy (mathmale):

Great. Now, one more. Can you evaluate \[50.28^{1/8}?\]

OpenStudy (anonymous):

1.631

OpenStudy (anonymous):

1.632

OpenStudy (mathmale):

Be sure to enclose the 1/8 with parentheses. Yes, 1.631 is fine. That's your r value. We used "fractional exponents" and "inverse functions" to obtain this result.

OpenStudy (mathmale):

Want to go on to another problem, or want to solve this same problem in a different way, using logarithms?

OpenStudy (anonymous):

No thanks for your help! :)

OpenStudy (mathmale):

My pleasure. Take care.

OpenStudy (anonymous):

You too :)